the median of the following set of numbers 12, 16, x+2, x+4, 32, 40 is 26 so find x?
Found 2 solutions by dfrazzetto, scott8148:
Answer by dfrazzetto(277)

6 numbers so the median is the average of the middle two
assuming a positive x, intuitively we see the numbers are already in order, as x+2 must be less than x+4
so:
(x+2+x+4)/2 = 26
2x+6= 52
2x= 46
x=23
Check, the average of 25 and 27 is indeed 26 √

Answer by scott8148(6628)

(12 + 16 + x+2 + x+4 + 32 + 40) / 6 = 26

2x + 106 = 26 * 6